# Hilbert’s fourth problem in the constant curvature setting

### v1

@@ 6.1. Basic properties
A necessity condition for the nonlinearity of $W^{1,p}_{0}(M,F,\operatorname{\mathfrak{m}})$ is given by the following result (cf. [20, Proposition 3.1]).
@@ 7. Asymmetric metrics on Euclidean spaces
The length $L_{F}(c)$ of a piecewise smooth curve $c$ and the induced distance $d_{F}$ are defined analogously to the Finsler case (see (2.29) and (2.30)). Then $d_{F}$ is an asymmetric metric on $\Omega$ satisfying the Busemann–Mayer theorem (cf. Bao et al. [1, pp. 160–161], Busemann–Mayer [9, Theorem 4.3, p. 186]), i.e.,
@@
for any smooth curve $\gamma:(-\epsilon,\epsilon)\rightarrow\Omega$ with $\gamma^{\prime}(0)=y\in T_{x}\Omega$. Moreover, it follows by [9, Theorem 4.2, p. 186] that $L_{F}(\gamma)=L_{d_{F}}(\gamma)$ for every piecewise smooth curve $\gamma:[0,1]\rightarrow\Omega$. Hence, we say that $F$ is projectively flat if $d_{F}$ is so, which coincides with the definition in the Finsler setting.
@@
Thus, $\gamma$ is an absolutely continuous curve in the sense of Rossi–Mielke–Savaré [29, p. 106]. By [29, Proposition 2.2], the limit
@@
A standard argument (cf. Burago–Burago–Ivanov [8, Theorem 2.7.4, Theorem 2.7.6]) then shows that
@@
Note that (7.9) remains valid for $\overline{F}$. Since $\overline{F}$ is quasi-regular in $y$, the Busemann–Mayer theorem [9, Theorem 4.3, p. 186] implies that
@@ Appendix A Complementary results for $\mathbf{K}\leq 0$
Alternatively, provided $\theta$ is not parallel to $y$, the positive definiteness of $\left([\bar{\psi}^{2}]_{y^{i}y^{j}}(y)\right)$ implies $\bar{\psi}_{y^{i}y^{j}}(y)\theta^{i}\theta^{j}>0$ (cf. [1, (1.2.9)]). Thus $\bar{F}^{2}]_{y^{i}y^{j}}\theta^{i}\theta^{j}>0$ follows by (A.43) and $\bar{\psi}(y)>0$, which completes the proof. ∎
@@ Appendix B Additional results for $\mathbf{K}=1$
Let $\sigma_{\alpha}:[0,2\pi]\rightarrow\mathbb{S}^{n}$, $\alpha\in\mathbb{N}$, be a sequence of great circles. By the Arzela–Ascoli theorem (cf. Burago et al. [8, Theorem 2.5.14]), there exists a subsequence $(\sigma_{\alpha_{k}})_{k}$ convergent uniformly to a closed curve $\sigma:[0,2\pi]\rightarrow\mathbb{S}^{n}$. Each $\sigma_{\alpha_{k}}$ is the intersection of $\mathbb{S}^{n}$ with a plane $\pi_{\alpha_{k}}\subset\mathbb{R}^{n+1}$ through the origin. Thus, for each $k$ and all $t\in[0,2\pi]$, we have
@@
In the remainder of this subsection, let $(x^{i})$ denote the standard Euclidean coordinates on $\mathbb{R}^{n}$ and $(\zeta^{j})=(\varphi,\theta^{1},\theta^{2},...,\theta^{n-1})$ the standard spherical coordinates on $\mathbb{S}^{n}$, where $\theta^{s}\in(0,\pi)$ for $s=1,\cdots,n-2$, $\theta^{n-1}\in[0,\pi)$, and $\varphi\in(-\frac{\pi}{2},\frac{3\pi}{2}]$. A direct (though computationally involved) calculation yields the following transformation formulas between the coframes $({\operatorname{d}}x^{i})$ and $({\operatorname{d}}\zeta^{j})$; see (LABEL:K=1spherical_x) for details.
@@ Appendix C Auxiliary properties of Sobolev spaces
For convenience, we denote by $\mathsf{Lip}(M)$ the collection of Lipschitz functions on $M$, and let $\mathsf{Lip}_{0}(M):=C_{0}(M)\cap\mathsf{Lip}(M)$ be the collection of Lipschitz functions with compact support. The following approximation (see [20, Proposition 3.3]) is useful in the sequel.

### v2

@@ 6.1. Basic properties
A necessity condition for the nonlinearity of $W^{1,p}_{0}(M,F,\operatorname{\mathfrak{m}})$ is given by the following result (cf. [34, Proposition 3.1]).
@@ 7. Asymmetric metrics on Euclidean spaces
The length $L_{F}(c)$ of a piecewise smooth curve $c$ and the induced distance $d_{F}$ are defined analogously to the Finsler case (see (2.29) and (2.30)). Then $d_{F}$ is an asymmetric metric on $\Omega$ satisfying the Busemann–Mayer theorem (cf. Bao et al. [4, pp. 160–161], Busemann–Mayer [17, Theorem 4.3, p. 186]), i.e.,
@@
for any smooth curve $\gamma:(-\epsilon,\epsilon)\rightarrow\Omega$ with $\gamma^{\prime}(0)=y\in T_{x}\Omega$. Moreover, it follows by [17, Theorem 4.2, p. 186] that $L_{F}(\gamma)=L_{d_{F}}(\gamma)$ for every piecewise smooth curve $\gamma:[0,1]\rightarrow\Omega$. Hence, we say that $F$ is projectively flat if $d_{F}$ is so, which coincides with the definition in the Finsler setting.
@@
Thus, $\gamma$ is an absolutely continuous curve in the sense of Rossi–Mielke–Savaré [43, p. 106]. By [43, Proposition 2.2], the limit
@@
A standard argument (cf. Burago–Burago–Ivanov [14, Theorem 2.7.4, Theorem 2.7.6]) then shows that
@@
Note that (7.9) remains valid for $\overline{F}$. Since $\overline{F}$ is quasi-regular in $y$, the Busemann–Mayer theorem [17, Theorem 4.3, p. 186] implies that
@@ Appendix A Complementary results for $\mathbf{K}\leq 0$
Alternatively, provided $\theta$ is not parallel to $y$, the positive definiteness of $\left([\bar{\psi}^{2}]_{y^{i}y^{j}}(y)\right)$ implies $\bar{\psi}_{y^{i}y^{j}}(y)\theta^{i}\theta^{j}>0$ (cf. [4, (1.2.9)]). Thus $\bar{F}^{2}]_{y^{i}y^{j}}\theta^{i}\theta^{j}>0$ follows by (A.43) and $\bar{\psi}(y)>0$, which completes the proof. ∎
@@ Appendix B Additional results for $\mathbf{K}=1$
Let $\sigma_{\alpha}:[0,2\pi]\rightarrow\mathbb{S}^{n}$, $\alpha\in\mathbb{N}$, be a sequence of great circles. By the Arzela–Ascoli theorem (cf. Burago et al. [14, Theorem 2.5.14]), there exists a subsequence $(\sigma_{\alpha_{k}})_{k}$ convergent uniformly to a closed curve $\sigma:[0,2\pi]\rightarrow\mathbb{S}^{n}$. Each $\sigma_{\alpha_{k}}$ is the intersection of $\mathbb{S}^{n}$ with a plane $\pi_{\alpha_{k}}\subset\mathbb{R}^{n+1}$ through the origin. Thus, for each $k$ and all $t\in[0,2\pi]$, we have
@@
In the remainder of this subsection, let $(x^{i})$ denote the standard Euclidean coordinates on $\mathbb{R}^{n}$ and $(\zeta^{j})=(\varphi,\theta^{1},\theta^{2},...,\theta^{n-1})$ the standard spherical coordinates on $\mathbb{S}^{n}$, where $\theta^{s}\in(0,\pi)$ for $s=1,\cdots,n-2$, $\theta^{n-1}\in[0,\pi)$, and $\varphi\in(-\frac{\pi}{2},\frac{3\pi}{2}]$. A direct (though computationally involved) calculation yields the following transformation formulas between the coframes $({\operatorname{d}}x^{i})$ and $({\operatorname{d}}\zeta^{j})$; see (5.17) for details.
@@ Appendix C Auxiliary properties of Sobolev spaces
For convenience, we denote by $\mathsf{Lip}(M)$ the collection of Lipschitz functions on $M$, and let $\mathsf{Lip}_{0}(M):=C_{0}(M)\cap\mathsf{Lip}(M)$ be the collection of Lipschitz functions with compact support. The following approximation (see [34, Proposition 3.3]) is useful in the sequel.